# A “Staircase” formula for the Chern-Schwartz-MacPherson cycle of a matroid

### v1

@@ 1 Introduction
The Chern-Schwartz-MacPherson (CSM) class generalizes the Chern class of the tangent bundle of a variety. In particular, it is defined even when the variety is singular or non-proper. One interpretation of the CSM class of a variety is in terms of a natural transformation between the functor of constructible functions on varieties and the Chow group functor, for more details, see [4]. In the case where $V$ is the complement of a hyperplane arrangement $\mathcal{A}$ in $\mathbb{P}^{d}$, a combinatorial formula for $c_{SM}(V)$ in terms of the characteristic polynomial of the underlying matroid was given separately by Aluffi in [5] and Huh in [15].
De Concini and Procesi defined the notion of a wonderful model of a hyperplane arrangement with respect to a building set in [8]. In this paper, we will use the term wonderful model of $\mathcal{A}$ to refer to the wonderful model with respect to the maximal building set. Later, Feichtner and Yuzhvinsky [10] define the Chow ring of a matroid following the ideas of [8] and the correspondence between the Chow ring of a nonsingular proper toric variety $X$ and Minkowski weights on its associated fan $\Sigma$ proven in [13]. In [1], Adiprasito, Huh and Katz bring these concepts together and prove that $A^{*}(M)$ and $MW_{*}(B(M))$ are dual to each other, in analogy to Poincaré duality for the toric case and the realizable case. This gave the necessary setting for López de Medrano, Rincón, and Shaw to define an analogous CSM cycle for matroids.
In [9], the authors define $\operatorname{csm}(M)$ for a matroid $M$ and prove in Theorem 3.1 that when $M$ is realizable over $\mathbb{C}$, its CSM cycle corresponds to the geometric CSM class $c_{SM}(\mathcal{W}\setminus\mathcal{A})$ of the complement of the arrangement in the wonderful model under the identification of $MW_{*}(B(M))\cong A_{*}(\mathcal{W})$. However, in the geometric setting $c_{SM}(\mathcal{W}\setminus\mathcal{A})$ should be Poincaré dual to some element $c^{SM}(\mathcal{W}\setminus\mathcal{A})\in A^{*}(\mathcal{W})$, since $\mathcal{W}$ is proper and nonsingular. In the combinatorial setting, the Minkowski weights on the Bergman fan of $M$ and the Chow ring of $M$ satisfy Poincaré duality, so there should be a class $c^{SM}(M)\in A^{*}(M)$ with $c^{SM}(M)\cap 1_{M}=\operatorname{csm}(M)$. In Section 3 we prove
@@
In [17], Fife and Rincón conjectured that the degree $k$ part of $c^{SM}(M)$ is given by the sum
@@
Theorem 3.1 proves this conjecture in the case of matroids realizable over $\mathbb{C}$ and Corollary 4.3 generalizes the answer to arbitrary matroids.
@@
The proof of this statement uses the contraction-deletion formula, the interpretation of the Chow ring in terms of piecewise polynomials [7] and induction on the number of elements of $M$, which completes the proof of Conjecture 5.2.4 in [17].
@@ 2 Background
For an in depth look at matroids, we reference [19].
@@
Let $\pi:\mathcal{W}\rightarrow\mathbb{P}^{d}$ denote this blow-up map. $\pi$ is an isomorphism on the complement of the arrangement $C(\mathcal{A})$. On the $\mathcal{W}$ there is a divisor corresponding to each flat of the matroid. All collections of these divisors intersect transversely if their flats are comparable, and are disjoint otherwise. See [8] for details.
@@
The following definitions of Chow ring and Minkowski weights of a fan $\Sigma$ are explored in detail in Chapter $5$ of [1]. We reproduce the basic definitions and facts for the reader:
@@
In the case where $\Sigma$ is the fan associated to a proper, nonsingular toric variety $X(\Sigma)$, $A^{*}(\Sigma)\cong A^{*}(X(\Sigma))$ and $A_{*}(X(\Sigma))\cong MW_{*}(\Sigma)$, see [13] for details.
@@ 3 A “Staircase” CSM Class Formula for Realizable Matroids
There exists a unique natural transformation $c_{*}:\mathcal{C}_{*}\to A_{*}$ such that for a proper, nonsingular variety $V$, $c_{*}(\mathbbm{1}_{V})=c(TV)\cap[V]$ where $\mathcal{C}_{*}:Var_{\mathbb{C}}\to Ab$ is the functor of constructible functions and $A_{*}:Var_{\mathbb{C}}\to Ab$ is the Chow group functor. See Theorem 1 in [16] for details.
@@
López de Medrano, Rincón, and Shaw defined an analogous CSM cycle for matroids, and proved that when a matroid is realizable over $\mathbb{C}$, its CSM cycle corresponds to the geometric CSM class $c_{SM}(\mathcal{W}\setminus\mathcal{A})$ of the complement of the arrangement in the wonderful model under the identification of $A^{*}(M)\cong A^{*}(\mathcal{W})$ (Theorem 3.1 in [9]).
A formula for the CSM cycle of a matroid in terms of the generators of the Chow Ring was conjectured by Fife, T. and Rincón, F in private communication, according to Conjecture 5.2.4 in [17].
@@
(Lemma 1.3 in [6]) Let $X\subseteq Y$ be a nonsingular varieties, and $\pi:\tilde{Y}\rightarrow Y$ the blow-up of Y along $X$, and $\tilde{X}$ the exceptional divisor. If $X$ is a complete intersection of hypersurfaces $Z_{1},\dots,Z_{d}$ meeting transversally in $Y$, then the Chern class of the tangent bundle of the blow up $\tilde{Y}$ is related to $c(T_{Y})$ by the formula
@@
(Theorem 1 in [3]) Let $X\subset Y$ be a subvariety of a nonsingular variety, with $X=D_{1}\cup\dots\cup D_{n}$ a normal crossings divisor with smooth components. The CSM class of the complement $Y\setminus X$ is
@@ 4.1 Contraction-deletion of $\operatorname{stair}$
For a given element $i\in E$ which is not a coloop, we obtain a projection $\delta:B(M)\to B(M\setminus i)$. By Proposition 2.22 in [18], $\delta$ is a tropical modification along some rational function $\phi$; therefore, $\delta$ induces a pullback $\delta^{*}:Z_{*}(B(M\setminus i)\to Z_{*}(B(M))$ and since $div_{B\setminus i}(\phi)=B(M/i)$, we also obtain a pullback $Z_{*}(B(M/i))\to Z_{*}(B(M\setminus i))$. These have corresponding ring/group homomorphisms in the context of the Chow rings, see Section 4.2 for details.
@@ 4.2 Piecewise polynomials and cocycles
There is an equivalent formulation of $A^{*}(\Sigma)$ in terms of so-called Courant functions on $\Sigma$. This approach defines $A^{*}(\Sigma)$ as a quotient of continuous piecewise polynomial functions on $\Sigma$ by linear functions from $N$ to $\mathbb{R}$, see Section 4 of [1] for more details.
This other formulation leads to the theory of tropical cocycles on $\Sigma$, introduced by François in [11] and elaborated upon by Gross and Shokrieh in [14], in analogy to the theory of cycles on $\Sigma$ from [2]. In particular, $Z_{*}(\Sigma)$, the tropical cycles on $\Sigma$, consists of weighted fans on subdivisions of $\Sigma$, while $C^{*}(\Sigma)$, the tropical cocycles on $\Sigma$, are continuous functions that are piecewise polynomial in a subdivision of $\Sigma$. In [14], the authors prove that, when given the corresponding intersection products, these $C^{*}(\Sigma)$-algebras are isomorphic. In particular, given a morphism of fans $f:\Sigma_{1}\to\Sigma_{2}$, a cocycle $h\in C^{*}(\Sigma_{2})$ and a cycle $X\in Z_{*}(\Sigma_{2})$, $f^{*}(h\cdot X)=f^{*}(h)\cdot f^{*}(X)$, where $f^{*}:Z_{*}(\Sigma_{2})\to Z_{*}(\Sigma_{1})$ is defined in Corollary 4.7 in [14], which agrees with Definition 2.16 in [18] in the case where $f$ is a tropical modification.
@@
Since $\hat{\delta}:\hat{B}(M)\to\hat{B}(M\setminus i)$ satisfies the conditions of Proposition 3.10 in [12], there exists a rational function $\phi$ such that $\delta$ is a tropical modification along $\phi$, and there is an isomorphism $\phi\cdot MW_{*}(\hat{B}(M\setminus i))\xrightarrow{\sim}MW_{*}(\hat{B}(M/i))$. The inverse of this isomorphism is given by $\hat{j}_{*}:MW_{*}(\hat{B}(M/i))\to MW_{*}(\hat{B}(M\setminus i))$.
@@
We proceed by induction. Note that for the loop matroid $M_{0}$, i.e. the matroid on one element whose only element is a loop, $\operatorname{csm}(M_{0})=0=0\cap 1_{M_{0}}=\operatorname{stair}(M_{0})\cap 1_{M_{0}}$ and for the isthmus $M_{1}$, i.e. the matroid on one element whose only element is independent, $\operatorname{csm}(M_{1})=1_{M_{1}}$. Now, assume that for all matroids $M^{\prime}$ whose ground set consists of $n$ elements $\operatorname{stair}(M^{\prime})\cap 1_{M^{\prime}}=\operatorname{csm}_{k}(M^{\prime})$ and let $M$ be a loopless matroid whose ground set consists of $n+1$ elements. If all elements of the ground set of $M$ are coloops, then $M$ is representable over $\mathbb{C}$ and thus, by Theorem 3.1, the conclusion follows. Now assume there exists $i\in M$ that is not a coloop. Then, by Theorem 4.1, $\operatorname{stair}(M)=\bar{\delta}^{*}\operatorname{stair}(M\setminus i)-\bar{\delta}^{*}\operatorname{stair}(M/i)$. Then, the ground sets of $M\setminus i$ and $M/i$ contain $n$ elements. Thus, by Lemma 4.2 and Theorem 5.4 in [9],

### v2

@@ 1 Introduction
The Chern-Schwartz-MacPherson (CSM) class generalizes the Chern class of the tangent bundle of a variety. In particular, it is defined even when the variety is singular or non-proper. One interpretation of the CSM class of a variety is in terms of a natural transformation between the functor of constructible functions on varieties and the Chow group functor, for more details and history, see this survey by Brasselet [Bra00]. In the case where $V$ is the complement of a hyperplane arrangement $\mathcal{A}$ in $\mathbb{P}^{d}$, a combinatorial formula for $c_{SM}(V)$ in terms of the characteristic polynomial of the underlying matroid was given separately by Aluffi in [Alu12] and Huh in [Huh13].
De Concini and Procesi defined the notion of a wonderful model of a hyperplane arrangement with respect to a building set in [DCP96]. In this paper, we will use the term wonderful model of $\mathcal{A}$ to refer to the wonderful model with respect to the maximal building set. Later, Feichtner and Yuzhvinsky [FY04] define the Chow ring of a matroid following the ideas of [DCP96] and the correspondence between the Chow ring of a nonsingular proper toric variety $X$ and Minkowski weights on its associated fan $\Sigma$ proven by Fulton and Strumfels in [FS97]. In [AHK18], Adiprasito, Huh and Katz bring these concepts together and prove that $A^{*}(M)$ and $MW_{*}(B(M))$ are dual to each other, in analogy to Poincaré duality for the toric case and the realizable case. This gave the necessary setting for López de Medrano, Rincón, and Shaw to define an analogous CSM cycle for matroids [LdMRS19].
In their paper [LdMRS19], López de Medrano et al. define $\operatorname{csm}(M)$ for a matroid $M$ and prove in Theorem 3.1 that when $M$ is realizable over $\mathbb{C}$, its CSM cycle corresponds to the geometric CSM class $c_{SM}(\mathcal{W}\setminus\mathcal{A})$ of the complement of the arrangement in the wonderful model under the identification of $MW_{*}(B(M))\cong A_{*}(\mathcal{W})$. They also generalize Aluffi and Huh’s statement about the relation between the CSM class and the characteristic polynomial to a general matroid. A separate proof of this relationship was provided by Ashraf and Backman in [AB22] using through an understanding of Tutte activities.
However, in the geometric setting $c_{SM}(\mathcal{W}\setminus\mathcal{A})$ should be Poincaré dual to some element $c^{SM}(\mathcal{W}\setminus\mathcal{A})\in A^{*}(\mathcal{W})$, since $\mathcal{W}$ is proper and nonsingular. In the combinatorial setting, the Minkowski weights on the Bergman fan of $M$ and the Chow ring of $M$ satisfy Poincaré duality, so there should be a class $c^{SM}(M)\in A^{*}(M)$ with $c^{SM}(M)\cap 1_{M}=\operatorname{csm}(M)$. In Section 3 we prove:
@@
In Mannino’s master thesis [Man22], Conjecture 5.2.4, attributed to a private communication between Tara Fife and Felipe Rincón, proposes that the degree $k$ part of $c^{SM}(M)$ is given by
@@
Theorem 3.1 proves this conjecture in the case of matroids realizable over $\mathbb{C}$ and Theorem 4.1 generalizes the answer to arbitrary matroids.
@@
The proof of this statement uses the contraction-deletion formula, the interpretation of the Chow ring in terms of piecewise polynomials [Bri96] and induction on the number of elements of $M$, which completes the proof of Conjecture 5.2.4 in [Man22].
Other work: In [Rau23], Rau provides a formula for the CSM cycle of a matroid $M$ in terms of the tropical polynomials cutting out the Bergman fan $1_{M}$, which is similar to Lemma 3.2 if we interpret the tropical polynomials as the sections whose zero sections are the components of the pullback of the hyperplane arrangement in the realizable case. In [BEST23], Berget, Eur, Spink and Tseng construct the framework of tautological classes of matroids, which generalizes the class of the universal sub and quotient bundle of the Grassmannian to an arbitrary matroid. Through this framework they obtain a formula for the inclusion of the group of Minkowski weights of a matroid $M$ into the Chow ring of the corresponding permutahedral variety and show that the Poincaré dual of $\operatorname{csm}(M)$ in the Chow ring of the permutahedral variety is $\sum_{k=0}^{\operatorname{rk}(M)-1}c_{k}(\mathcal{S_{M}})$. Although our proof of Theorem 4.1 does not use this framework, our result gives the Chern roots of $[\mathcal{S}_{M}]$ as $\{-\sum_{\begin{subarray}{c}F\in\mathcal{L}\\ \operatorname{rk}(F)\geq r\end{subarray}}x_{F}|1\leq r\leq\operatorname{rk}(M)\}$ in $A^{*}(M)$, via appendix III in [BEST23].
@@ 2 Background
For an in depth look at matroids, we reference [Whi86].
@@
Let $\pi:\mathcal{W}\rightarrow\mathbb{P}^{d}$ denote this blow-up map. $\pi$ is an isomorphism on the complement of the arrangement $C(\mathcal{A})$. On the $\mathcal{W}$ there is a divisor corresponding to each flat of the matroid. All collections of these divisors intersect transversely if their flats are comparable, and are disjoint otherwise. See [DCP96] for details.
@@
The following definitions of Chow ring and Minkowski weights of a fan $\Sigma$ are explored in detail in Chapter $5$ of [AHK18]. We reproduce the basic definitions and facts for the reader:
@@
In the case where $\Sigma$ is the fan associated to a proper, nonsingular toric variety $X(\Sigma)$, $A^{*}(\Sigma)\cong A^{*}(X(\Sigma))$ and $A_{*}(X(\Sigma))\cong MW_{*}(\Sigma)$, see [FS97] for details.
@@ 3 A “Staircase” CSM Class Formula for Realizable Matroids
There exists a unique natural transformation $c_{*}:\mathcal{C}_{*}\to A_{*}$ such that for a proper, nonsingular variety $V$, $c_{*}(\mathbbm{1}_{V})=c(TV)\cap[V]$ where $\mathcal{C}_{*}:Var_{\mathbb{C}}\to Ab$ is the functor of constructible functions and $A_{*}:Var_{\mathbb{C}}\to Ab$ is the Chow group functor. See Theorem 1 in [Mac74] for details.
@@
López de Medrano, Rincón, and Shaw defined an analogous CSM cycle for matroids, and proved that when a matroid is realizable over $\mathbb{C}$, its CSM cycle corresponds to the geometric CSM class $c_{SM}(\mathcal{W}\setminus\mathcal{A})$ of the complement of the arrangement in the wonderful model under the identification of $A^{*}(M)\cong A^{*}(\mathcal{W})$ (Theorem 3.1 in [LdMRS19]).
A formula for the CSM cycle of a matroid in terms of the generators of the Chow ring was conjectured by Fife, T. and Rincón, F in private communication, according to Conjecture 5.2.4 in [Man22].
@@
(Lemma 1.3 in [Auf09]) Let $X\subseteq Y$ be a nonsingular varieties, and $\pi:\tilde{Y}\rightarrow Y$ the blow-up of Y along $X$, and $\tilde{X}$ the exceptional divisor. If $X$ is a complete intersection of hypersurfaces $Z_{1},\dots,Z_{d}$ meeting transversally in $Y$, then the Chern class of the tangent bundle of the blow up $\tilde{Y}$ is related to $c(T_{Y})$ by the formula
@@
(Theorem 1 in [Alu99]) Let $X\subset Y$ be a subvariety of a nonsingular variety, with $X=D_{1}\cup\dots\cup D_{n}$ a normal crossings divisor with smooth components. The CSM class of the complement $Y\setminus X$ is
@@ 4.1 Contraction-deletion of $\operatorname{stair}$
For a given element $i\in E$ which is not a coloop, we obtain a projection $\delta:B(M)\to B(M\setminus i)$. By Proposition 2.22 in [Sha13], $\delta$ is a tropical modification along some rational function $\phi$; therefore, $\delta$ induces a pullback $\delta^{*}:Z_{*}(B(M\setminus i)\to Z_{*}(B(M))$ and since $div_{B\setminus i}(\phi)=B(M/i)$, we also obtain a pullback $Z_{*}(B(M/i))\to Z_{*}(B(M\setminus i))$. These have corresponding ring/group homomorphisms in the context of the Chow rings, see Section 4.2 for details.
@@ 4.2 Piecewise polynomials and cocycles
There is an equivalent formulation of $A^{*}(\Sigma)$ in terms of so-called Courant functions on $\Sigma$. This approach defines $A^{*}(\Sigma)$ as a quotient of continuous piecewise polynomial functions on $\Sigma$ by linear functions from $N$ to $\mathbb{R}$, see Section 4 of [AHK18] for more details.
This other formulation leads to the theory of tropical cocycles on $\Sigma$, introduced by François in [Fra13] and elaborated upon by Gross and Shokrieh in [GS19], in analogy to Allermann and Rau’s theory of cycles on $\Sigma$ from [AR09]. In particular, $Z_{*}(\Sigma)$, the tropical cycles on $\Sigma$, consists of weighted fans on subdivisions of $\Sigma$, while $C^{*}(\Sigma)$, the tropical cocycles on $\Sigma$, are continuous functions that are piecewise polynomial in a subdivision of $\Sigma$. In [GS19], the authors prove that, when given the corresponding intersection products, these $C^{*}(\Sigma)$-algebras are isomorphic. In particular, given a morphism of fans $f:\Sigma_{1}\to\Sigma_{2}$, a cocycle $h\in C^{*}(\Sigma_{2})$ and a cycle $X\in Z_{*}(\Sigma_{2})$, $f^{*}(h\cdot X)=f^{*}(h)\cdot f^{*}(X)$, where $f^{*}:Z_{*}(\Sigma_{2})\to Z_{*}(\Sigma_{1})$ is defined in Corollary 4.7 in [GS19], which agrees with Definition 2.16 in [Sha13] in the case where $f$ is a tropical modification.
@@
Since $\hat{\delta}:\hat{B}(M)\to\hat{B}(M\setminus i)$ satisfies the conditions of Proposition 3.10 in [FR10], there exists a rational function $\phi$ such that $\delta$ is a tropical modification along $\phi$, and there is an isomorphism $\phi\cdot MW_{*}(\hat{B}(M\setminus i))\xrightarrow{\sim}MW_{*}(\hat{B}(M/i))$. The inverse of this isomorphism is given by $\hat{j}_{*}:MW_{*}(\hat{B}(M/i))\to MW_{*}(\hat{B}(M\setminus i))$.
@@
We proceed by induction. Note that for the loop matroid $M_{0}$, i.e. the matroid on one element whose only element is a loop, $\operatorname{csm}(M_{0})=0=0\cap 1_{M_{0}}=\operatorname{stair}(M_{0})\cap 1_{M_{0}}$ and for the isthmus $M_{1}$, i.e. the matroid on one element whose only element is independent, $\operatorname{csm}(M_{1})=1_{M_{1}}$. Now, assume that for all matroids $M^{\prime}$ whose ground set consists of $n$ elements $\operatorname{stair}(M^{\prime})\cap 1_{M^{\prime}}=\operatorname{csm}_{k}(M^{\prime})$ and let $M$ be a loopless matroid whose ground set consists of $n+1$ elements. If all elements of the ground set of $M$ are coloops, then $M$ is representable over $\mathbb{C}$ and thus, by Theorem 3.1, the conclusion follows. Now assume there exists $i\in M$ that is not a coloop. Then, by Theorem 4.1, $\operatorname{stair}(M)=\bar{\delta}^{*}\operatorname{stair}(M\setminus i)-\bar{\delta}^{*}\operatorname{stair}(M/i)$. Then, the ground sets of $M\setminus i$ and $M/i$ contain $n$ elements. Thus, by Lemma 4.2 and Theorem 5.4 in [LdMRS19],
@@
In the context of tautological classes of matroids:
Remark 4.1.
For an arbitrary matroid $M$, the class $\sum_{i=0}^{\operatorname{rk}(M)-1}c_{\operatorname{rk}(M)-1-k}(\mathcal{S_{M}})$ equals the staircase of $M$. In particular, the Chern roots of $S_{M}$ are
$\alpha_{r}=\sum_{\begin{subarray}{c}F\in\mathcal{L}(M)\\ \operatorname{rk}(F)\geq r\end{subarray}}x_{F}.$
Note that this is different, yet implied by Appendix III in [BEST23], as the use of modular filters is not needed and the computation happens in $A^{*}(M)$.